find the value of x.
3x + 15
4x + 12

Answer by cloudstrifeo0o(5)

This is how I was taught:
3x + 15 | 4x + 12
Set the equation to 0.
3x + 15 = 0 | 4x + 12 = 0
-15 -15 | -12 -12
Subtract by 15. | Subtract by 12
3x = -15 | 4x = -12
/3 /3 | /4 /4
Divide by 3 | Divide by 4
x=-5 | x=-3
I could be wrong, but I'm pretty sure it's right.